OpenStudy (anonymous):

If t varies as v, and t=2 4/7 when v= 13/14 find v when t= 2 1/4

OpenStudy (jdoe0001):

\(\bf t = (n)v\qquad t = 2\frac{4}{7}\qquad v=2\frac{1}{4}\\ \quad \\ t=(n)v\implies 2\frac{4}{7}=(n)2\frac{1}{4}\implies \cfrac{2\frac{4}{7}}{2\frac{1}{4}}=n\)

OpenStudy (jdoe0001):

hmmm I put the wrong... fraction..... one sec

OpenStudy (jdoe0001):

\(\bf t = (n)v\qquad t = 2\frac{4}{7}\qquad v=\frac{13}{14}\\ \quad \\ t=(n)v\implies 2\frac{4}{7}=(n)\frac{13}{14}\implies \cfrac{2\frac{4}{7}}{\frac{13}{14}}=n\)

OpenStudy (anonymous):

the answer choices are 2106/392 13/16 324/52

OpenStudy (jdoe0001):

\(\bf t = (n)v\implies \cfrac{t}{(n)}=v\qquad \\ \quad \\ when \qquad t=2\frac{1}{4}\qquad \cfrac{t}{(n)}=v\implies \cfrac{2\frac{1}{4}}{(n)}=v\)

OpenStudy (loser66):

:)